January 8th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • recovering from power outage - taking more cosmics - might be on bad
      • will try pulser
    • Bigite shower
      • issue with HV, fixed after reboot
      • tedbbdaq froze, and could not log in, need to investigate - no warning in log files - might be running out of memory or issue with NFS mounts - seems issues with local terminal issue not seen with GEM data taking
      • FADC crate next on list - give hostname VTP to b
      • atedf3 had to be powered on after power outage
    • BigBite hodoscope
      • Detector fully reassembled, can see signals
      • Ralph tried running CODA
    • CDet
      • nothing new for now - might be ready to take data
      • VXS shipped - VETROC
    • GEM
      • GEM UVA
        • DAQ for all layers tested - 3 layers installed - everything looks OK - will take cosmics next week
        • new cables being made for LV
      • GEM INFN
      • GEM TEDF
        • setup working with a few APVs using internal pulser : 5 APV connected but only see 4 APVs most of the time, Thir will add cables to have same latency has in other setup
        • will get additionnal APV backpplane
        • SSP decoder from Xinzhan
        • ready to try new firmware but will try after system stable
        • need setup L1A for MPD
    • GRINCH : CODA3 running with updated drivers - Bradley can start testing - new VETROC config using config file